OpenStudy (fools101):

Exponential Functions The population of a mountain town in Colorado was 4600 in 1980 and in 1990 grew to 5000. Write an exponential equation that models this information

OpenStudy (anonymous):

is this supposed to be solved by first order differential equations. Ie any more info for the question?

OpenStudy (anonymous):

as in dy/dt = ky? cuz that gives y = Ce^(kt)

OpenStudy (anonymous):

where y0 = C = 4600 and k= (lny)/t or k = ln(5000)/10 so basically its described as y = 4600e^((ln(5000)/10)*t so thats your equation you are looking for i assume

OpenStudy (anonymous):

answer your question? :D

OpenStudy (fools101):

Yes i got it now :D thank u Zucchini that helps a lot!!!!

OpenStudy (fools101):

Thank u!

OpenStudy (anonymous):

np ;)

OpenStudy (fools101):

^ ^!!!

OpenStudy (anonymous):

ODEs are always difficult ;) gd luck with your hw

OpenStudy (fools101):

i know i'm starting to see that lol and thank u :)!